# supplement_arXiv

### v2

@@ 5 Simulations
In the following two sections, we investigate the robustness of Algorithm 1 against data sets where a fixed proportion $c$ of the data has been replaced by outliers and compare it to the non-robust SFPCA method.
@@ 5.1 Simulation 1
In the first simulation, in each of $100$ repetitions, we consider a data set of size $n=200$ consisting of $n-\lfloor cn\rfloor$ regular- and $\lfloor cn\rfloor$ contaminated curves. Each curve is estimated using kernel density estimation (KDE) on a generated sample and evaluated at an equidistant grid of $p=50$ time points. The construction of a regular/outlying sample is as follow:
@@
where $Y_{i}\sim N(0,1),~{}\tilde{Y}_{i}\sim U([q1;q2]\cup[-q2;-q1]),N=250$ and $q_{1},q_{2}$ being the $0.1\%$ and $0.5\%$ quantile of the standard normal distribution, respectively. KDE was calculated using the R-package kde1d (Nagler, 2024).
@@
To measure the performance of RDPCA, we consider two statistics: the integrated square error (ISE), approximating the norm of the difference between estimates of the true and sub-sampled covariances, and the mean cosine between the first five pairs of eigenfunctions based on the same estimates. For more details, see Supplement Section LABEL:supp:simulations.
During the analysis, we consider contamination rates $c\in\{0,0.05,0.1,0.2\}$, automatically select the regularization parameter $\alpha$ according to the procedure discussed in Section 4.2. Furthermore, we fix $h=\lceil 0.75n\rceil$ and $k=1$. The results w.r.t. contamination rate $c$ are shown in Figure 2, where the mean value over 100 repetitions +/+/-+ / - one standard error are depicted. Both the ISE and the mean cosine suggest that RDPCA significantly outperforms SFPCA in the presence of outliers.
@@
Figure 2: Integrated square error (left) and mean cosine (right) between first five pairs of estimated and true eigenfunctions. Displayed are the mean +/+/-+ / - one standard error of each measurement over all iterations at each level of the corresponding parameter configuration. All measurements are calculated based on the covariance estimated by RDPCA (solid) and SFPCA (dashed) and respectively compared with the estimated true covariance.
In Figure 3, the robust correlation function nicely aligns with the correlation function based on a Monte Carlo estimate of the true underlying model. On the contrary, the non-robust correlation reveals significantly different structure. Below, in Figure 4 we compare the squared RDMD based on these two sample estimates and the squared robust $\alpha$-Mahalanobis (Berrendero et al., 2020) distance, ignoring the compositional structure of the densities. The dashed lines indicate the $95\%$ quantile of the limiting distribution of these distances assuming Gaussianity; see Corollary 4.1. These values are estimated by Monte Carlo simulation. Then, a density is considered outlying if its squared Mahalanobis distance exceeds this cutoff. The circles represent the true regular observations, whereas the triangles correspond to the outliers. According to the cutoff values, only RDPCA can determine outliers in a meaningful way. In the supplement, we added some additional results related to this comparison.
Figure 3: True (left), RDPCA with $\alpha=0.48$ (middle), and SFPCA (right) correlation function, for the example shown in Figure 1.
Figure 4: Distance-distance plot of squared robust vs. non-robust RDMD (left) and robust $\alpha$-Mahalanobis (Berrendero et al., 2020) (right) distances for example of Figure 1. The regularization parameter is $\alpha=0.48$. Dashed lines indicate the corresponding cutoff values, under Gaussianity. Circles and triangles correspond to true regular and outlying observations, respectively.
@@ 5.2 Simulation 2
In the second simulation, we generate data sets of size $n=200$ according to
$\mathrm{clr}(X_{\mathrm{reg}}(t))=$ $~{}s_{1}\xi_{1}+s_{2}\xi_{2}+s_{3}\xi_{3}+s_{4}\xi_{4},$
$\mathrm{clr}(X_{\mathrm{out}}(t))=$ $~{}s_{1}\xi_{1}+s_{2}\xi_{2}+s_{3}\xi_{3}+s_{4}\xi_{4}+s_{5}\xi_{5},$
[…]
We either consider scores that independently follow a normal distribution, i.e. $s_{i}\sim N(0,\lambda_{i})$ or jointly a multivariate t-distribution, i.e $(s_{1},s_{2},s_{3},s_{4})\sim t_{5}(\boldsymbol{0},\mathrm{diag}(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}))$ and $(s_{1},s_{2},s_{3},s_{4},s_{5})\sim t_{5}(\boldsymbol{0},\mathrm{diag}(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4},\lambda_{5}))$. In both cases we set $(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4},\lambda_{5})=(2,1,1/2,1/4,4)$. The resulting process is then evaluated at an equidistant grid of $p=100$ time points. Figure 5 shows an example of the proposed model. For more details on the extension and the analysis of a second model, we refer to the supplement.
Figure 5: Visualization of curves for normal (left) and multivariate t-distributed (right) scores. Solid curves represent the main processes, dashed ones indicate outliers.
Each setting is again replicated 100 times for $c=\{0,0.05,0.1,0.2\}$ and analyzed using automatically selected $\alpha$, $h=\lceil 0.75n\rceil$, $k=1$. The accuracy of RDPCA is evaluated by the ISE of the difference between the estimated and the true covariance, as well as the mean cosine between the first four pairs of eigenfunctions based on the estimated and the true covariance. The corresponding results are shown in Figure 6.
Figure 6: Integrated square error (left) and cosine between first four pairs of estimated and true eigenfunctions (right). Displayed are the mean +/+/-+ / - one standard error of each measurement over all iterations at each level of the corresponding parameter configuration. All measurements are calculated based on the covariance estimated by RDPCA (solid) and SFPCA (dashed) and respectively compared with a true covariance function. The notation “normal" and “t" refer to the distribution of the scores in the underlying model as described at the beginning of this section.
In the final analysis, we contrast the estimated ratio of explained variance between robust and non-robust approaches to the true proportions. The results presented in Figure 7 show that the robust methodology is able to accurately estimate the true ratios.
@@ 6.1 EPXMA spectra
The considered dataset consists of electron probe X-ray microanalysis (EPXMA) spectra of various glass vessel types. This type of analysis is commonly used to determine the chemical composition of glasses by measuring X-rays at different wavelength emitted from a sample after being hit by a beam of electrons. The resulting intensities are inherently relative. Furthermore, due to the mutual influence chemical elements have on each other during this process, it is reasonable to view EPXMA spectra as relative functional data.
The dataset contains 180 such observations divided into four distinct groups corresponding to different glass vessel types. The largest group, referred to as “sodic", comprises 145 spectra. The remaining groups, designated as “potassic", “potasso-calcic", and “calcic" are represented by $15,10,$ and $10$ spectra, respectively. It is important to note that the latter 38 spectra within the sodic group were acquired under varying experimental conditions. Each observation is characterized by measurements across 1920 wavelengths. For the purposes of this analysis, we focus on the initial 750 wavelengths as these encompass the primary variability observed in the data; for more details see Lemberge et al. (2000). This results in $n=180,\,p=750$.
The left plot of Figure 8 shows the trajectories of all clr transformed observations colored according to their respective glass vessel type. For the analysis, we consider the first 107 observations from the sodic group as regular observations. We select $\alpha=0.09$ as in Section 4.2. Additionally, we have $h=\lceil 0.5n\rceil$ and $k=1$.
@@
Consequently, the structure of the covariance or correlation function can also be significantly influenced in the presence of outlying curves. This is showcased in Figure 9 where robust and non-robust correlations are compared. Between wavelength 100 to 250 and around 400, the outliers exhibit a different structure. Especially at these parts of the domain, the non-robust correlation shows its bias towards the outlying curves.
@@ 6.2 Fertility data
The second data set considered in this study focuses on age-specific fertility rates (ASFR) across different countries. ASFR represents the number of women giving birth per $1,000$ within specific age groups, providing a detailed view of how fertility changes with age. The resulting curves are analyzed as density data because we are interested in the underlying distribution across the different age groups rather then the magnitude of it. We examine data from the Human Fertility Database, freely available at
https://www.humanfertility.org/Home/Index, focusing on the years $1956$, $1990$, and $2019$, to gain insight in the evolution fertility dynamics.
In the left column of Figure 10 we display the fertility densities. Outlying countries according to Algorithm 1 are represented by dashed curves. During the early stages of this period, peak fertility was primarily achieved by young women in their early 20s. Large families were common, reflected in a rather slow decrease in the majority of the fertility curves. During the following decades, mostly influenced by the introduction of reliable contraceptives and a change in the mindset about life quality, later first-pregnancies became more and more common. Throughout Europe, this phenomenon started earlier in the West than in the East, as can be seen by the fertility curves corresponding to the year $1990$ Since then, the peak fertility has steadily increased and sits now at around 30 years. For more details we refer to the insightful article by Beaujouan (2020) and references therein.
While the variability in the Bayes space seems to be present within the peaks, the upper row of Figure 10 shows that the main variability within the clr space is significantly higher in the tails of the distribution; see also (Hron et al., 2016).
@@
In Figure 11 the first two principal functions are displayed. In the brackets, we included the corresponding percentage of explained variance. As assessed before, for each year, the first component mainly captures the variability in the tails of the densities. Here, the principal functions focus on the left tail and the following ages until the early 20s. The second principal function contains significantly less information mainly explaining the right tail and the peak fertility.
@@
Following the construction of robust principal components, we look at the corresponding estimated principal component scores $\hat{s}_{j}$ associated with each country. In Figure 12 the first two robust scores for each country are visualized. For 1956 the outliers, as indicated by the triangles, all lay at the border of the data cloud. For Japan, both scores are exceedingly high, indicating not only relatively low fertility in younger age groups but also proportionally very high peak fertility. For the outliers of 1990 and 2019, we obtain a similar picture dominated by mostly extreme scores. For 1990 this mainly reflects countries that have a very early peak fertility. Similarly, in 2019 these extreme values explain high fertility rates at younger ages as well as late peak fertility.
Figure 12: Robust principal component scores for the first and second eigenfunction for the years $1956,1990,2019$ (from left to right). Circles correspond to regular observations while the triangles indicate outlying countries.
@@ 7 Conclusions and discussion
Despite the attention that the Bayes space methodology for density data analysis has attracted over the last decades, little focus has been paid to robust frameworks necessary for meaningful analysis in the presence of anomalies. Therefore, this paper introduces robust density PCA (RDPCA) as a methodology to robustly estimate the covariance operator and functional principal components of densely sampled univariate density data. Within this newly proposed method, we extend the notion of a functional Mahalanobis distance (Berrendero et al., 2020; Galeano et al., 2015) to the Bayes space. The resulting regularized Mahalanobis distance (RDMD) leads to a robust covariance estimation based on a subsample of those curves that best fit the underlying distribution. Based on this estimate, robust PCs for density function can be determined as summarized in Algorithm 1. During simulations, we demonstrated the efficiency of RDPCA as well as the advantages it has against a non-robust approach in the presence of outliers. Furthermore, if outliers arise within the tails of the densities, the clr transformation can amplify these deviations. Especially in these scenarios, we observed that common robust procedures that do not respect the nature of the data would fail.
@@
One of our initial assumptions was that the data had been observed at a dense grid. In the case where the data is only available at a sparse grid, further extensions would require smoothing by an appropriate basis, e.g., CB-splines (Machalová et al., 2021) specifically developed for Bayes spaces. Naturally, the next step from univariate data would be to consider multivariate densities. During multivariate functional data analysis, each observation contains the recording of several “functional" variables. In this setting, not only the covariance between different time points is considered, but also the relation between individual variables. If the second dimension is also continuously observed, the observations are so-called random surfaces. As these statistical fields show increasingly practical relevance (Berrendero et al. (2011); Górecki et al. (2018); Dai and Genton (2018); Masak and Panaretos (2023) and references therein), expanding RDPCA to these areas seems worthwhile. Calculating the RDMD involved regularizing by a suitable operator that smooths out unwanted noise components while keeping the relevant signal within the (uncontaminated) data unaffected. Another class of meaningful operators that utilize the functional nature of the data are differential operators often used in Tikhonov regularization. One has to keep in mind that these operations have yet to be defined for Bayes spaces. Next to PCA, the regularized Mahalanobis distance could be used for concepts like linear or quadratic discriminant analysis for the classification density data. As these methods rely on similarity measures involving several covariance operators, the robust classification of densities based on the RDMD would be suitable.
@@
The supplement contains 3 sections. Within Section LABEL:supp:simulations, we provide additional simulation results corresponding to Section 5.1 and 5.2 of this work. In Section LABEL:supp:realdata, further results of the analysis of the fertility data in Section 6.2 are supplied. Section LABEL:supp:proofs finishes with the proofs.

### v3

@@ 5 Simulations
Within two simulation studies, we evaluate the accuracy of the proposed robust approach in estimation of the covariance and principal components under contamination with a fixed proportion $c\in\{0,0.05,0.1,0.2\}$ of the data is replaced by outliers. Outliers are introduced via Huber’s $\varepsilon$-contamination model Huber (1964) as $X\sim(1-c)F_{reg}+cF_{out}$, where $F_{reg}$, $F_{out}$ denote the distribution of the regular and outlying observations, respectively.
To mitigate the effect of the data processing on the results, in the simulations we use the raw data sampled on a dense grid and robustly estimate the covariance and PCs using Algorithm LABEL:alg:alg1 from the supplement, with the regularization parameter $\alpha$ automatically chosen according to the procedure discussed in Section 4.2. Furthermore, we fix $h=\lfloor 0.75n\rfloor$ and $k=1$.
We assess both the accuracy and robustness of the covariance operator and principal component estimation, and compare the results with those of the non-robust SFPCA method. To measure the performance of RDPCA, we consider two statistics: the integrated square error (ISE), approximating the norm of the difference between estimates of the true and sub-sampled covariances, and the mean cosine between the first five pairs of eigenfunctions based on the same estimates. For more details, see Supplement Section LABEL:supp:simulations. For completeness, we also examine the performance of several outlier detection methods applied after the clr transformation, as commonly done in the functional data analysis literature (Arribas-Gil and Romo, 2014; Berrendero et al., 2020; Oguamalam et al., 2024). Detailed results are provided in Supplementary Section LABEL:sec:add_simul_results_supp.
@@ 5.1 Simulation 1
In the first simulation, in each of $100$ replications, we consider a data set of size $n=500$ consisting of $n-\lfloor cn\rfloor$ regular- and $\lfloor cn\rfloor$ contaminated curves, $c\in\{0,\,0.05,\,0.1,\,0.2\}$. Each curve is estimated using kernel density estimation (KDE) on a generated sample and evaluated at an equidistant grid of $p=50$ time points. The construction of a regular/outlying sample is as follows:
@@
where $Y_{i}\sim N(0,1),~\tilde{Y}_{i}\sim U([q_{1};q_{2}]\cup[-q_{2};-q_{1}]),N=250$ and $q_{1},q_{2}$ being the $0.1\%$ and $0.5\%$ quantile of the standard normal distribution, respectively. KDE was calculated using the R-package kde1d (Nagler, 2024).
@@
Figure 2 depicts the accuracy of the estimation of principal components w.r.t. the contamination rate $c$. Both the ISE and the mean cosine between the first five pairs of principal functions suggest that RDPCA significantly outperforms SFPCA in the presence of outliers. Interestingly, the right panel of Figure 2 indicates that RDPCA outperforms SFPCA in estimating the principal functions even in the non-contaminated setting.
@@
Figure 2: Mean ($\pm$ standard error) ISE between the estimated and true covariances (left) and cosine similarity between the first five pairs of estimated and true eigenfunctions (right), obtained using RDPCA (solid) and SFPCA (dashed).
Figure 3 illustrates how the robust correlation function nicely aligns with the correlation function based on a Monte Carlo estimate of the true underlying model. On the contrary, the non-robust correlation reveals a significantly different structure. In Figure 4 we compare the squared RDMD computed using RDPCA to the squared RDMD obtained using SPCA and the squared robust $\alpha$-Mahalanobis (Berrendero et al., 2020) distance, ignoring the compositional structure of the densities. The dashed lines indicate the $95\%$ quantile of the limiting distribution of these distances assuming Gaussianity and estimated by Monte Carlo simulation; see Corollary 4.1. A density is considered outlying if its squared Mahalanobis distance exceeds this cutoff. The circles represent the true regular observations, whereas the triangles correspond to the outliers. According to the cutoff values (dashed lines), only RDPCA can determine outliers in a meaningful way. In the supplement, we added some additional results related to this comparison.
Figure 3: True (left), RDPCA with $\alpha=0.32$ (middle), and SFPCA (right) correlation function, for the example shown in Figure 1.
Figure 4: Distance-distance plot of squared robust vs. non-robust RDMD (left) and robust $\alpha$-Mahalanobis (Berrendero et al., 2020) (right) distances for example of Figure 1. The regularization parameter is $\alpha=0.32$. Dashed lines indicate the corresponding cutoff values under Gaussianity. Circles and triangles correspond to true regular and outlying observations, respectively.
@@ 5.2 Simulation 2
In the second simulation, we generate data sets of size $n=500$ according to
$\mathrm{clr}(X_{\mathrm{reg}}(t))=$ $~s_{1}\xi_{1}+s_{2}\xi_{2}+s_{3}\xi_{3}+s_{4}\xi_{4},$ (14)
$\mathrm{clr}(X_{\mathrm{out}}(t))=$ $~s_{1}\xi_{1}+s_{2}\xi_{2}+s_{3}\xi_{3}+s_{4}\xi_{4}+s_{5}\xi_{5},$ (15)
where $\xi_{1}=\sqrt{2}\mathrm{sin}(2\pi t),\xi_{2}=\sqrt{2}\mathrm{cos}(2\pi t),\xi_{3}=\sqrt{2}\mathrm{sin}(4\pi t),\xi_{4}=\sqrt{2}\mathrm{cos}(4\pi t)$ are the eigenfunctions of the covariance operator of the clr transformed densities. The scores $s_{1},s_{2},s_{3},s_{4}$ are drawn, in one setting, from a normal distribution, $s_{i}\sim N(0,\lambda_{i})$ or, in another setting, from a multivariate t-distribution; $(s_{1},s_{2},s_{3},s_{4})\sim t_{5}(\boldsymbol{0},\mathrm{diag}(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}))$ and $(s_{1},s_{2},s_{3},s_{4},s_{5})\sim t_{5}(\boldsymbol{0},\mathrm{diag}(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4},\lambda_{5}))$. In both cases we set $(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4},\lambda_{5})=(2,1,1/2,1/4,4)$. The resulting process is then evaluated at an equidistant grid of $p=100$ time points. Visualization of the data is given in the supplement in Figure LABEL:fig:dataClrSim2Mod12.
Each setting is replicated 100 times, and the accuracy of RDPCA is evaluated by the ISE of the difference between the estimated and the true covariance, as well as the mean cosine between the first four pairs of eigenfunctions based on the estimated and the true covariance. The results are shown in Figure 5, and again demonstrate that the robust procedure outperforms SFPCA in the presence of outliers. Interestingly, for $t$-distributed scores, even in the setting with no outliers, the robust procedure yields better results.
Figure 5: Mean ($\pm$ standard error) ISE between the estimated and true covariances (left) and cosine similarity between the first five pairs of estimated and true eigenfunctions (right), obtained using RDPCA (solid) and SFPCA (dashed). The notation “normal" and “t" refer to the distribution of the scores in the underlying model as given at the beginning of the section.
Finally, we compare the estimated ratios of explained variance from the robust and non-robust methods to the true values. As shown in Figure 6, the robust method provides estimates that closely match the true ratios.
@@ 6 Real-data example
To illustrate the practical performance of RDPCA, we apply it to two real-data examples. The computations were carried out on a machine equipped with an AMD Ryzen 9 5900HX core with Radeon Graphics (16 cores, 3.30 GHz) and 32 GB of RAM, running Windows 11 Pro. The implementation was written in R (version 4.4.2). Calculations connected to RDPCA were run using 15 parallel cores and required approximately 5 and 1 minute(s) for the first and second example, respectively.
In addition to robust PCA, during the examples, we also examine atypical curves according to Algorithm 1. For this, we estimate the $95\%$-quantile of the limiting distribution of the RDMD under the assumption of Gaussianity; see Corollary 4.1. A density is then considered outlying if its squared RDMD exceeds this cutoff.
@@ 6.1 EPXMA spectra
The considered dataset consists of electron probe X-ray microanalysis (EPXMA) spectra of various glass vessel types. This type of analysis is commonly used to determine the chemical composition of glasses by measuring X-rays at different wavelengths emitted from a sample after being hit by a beam of electrons. As is commonly done in the analysis of spectral data, we analyze EPXMA spectra as relative functional data.
The dataset contains 180 such observations divided into four distinct groups corresponding to different glass vessel types. The largest group, referred to as “sodic", comprises 145 spectra. The remaining groups, designated as “potassic", “potasso-calcic", and “calcic", are represented by $15,10,$ and $10$ spectra, respectively. It is important to note that the latter 38 spectra within the sodic group were acquired under varying experimental conditions. Each observation is characterized by measurements across 1920 wavelengths. For the purposes of this analysis, we focus on wavelengths 51–750 as these encompass the primary variability observed in the data and to avoid measurement artifacts; for more details, see Lemberge et al. (2000). This results in n=180,p=700n=180,\,p=700.
We analyze the data using Algorithm 1, where the data is expressed using $31$ compositional splines. The left plot of Figure 7 shows the smoothed trajectories of the clr transformed observations colored according to their respective glass vessel type. For the analysis, we consider the first 107 observations from the sodic group as regular observations. We use $h=\lceil 0.5n\rceil$ and $k=4$, and automatically select $\alpha=0.03$; see Section 4.2.
@@
Figure 8 showcases the consequent influence onto the covariance, where we compare smoothed robust to non-robust correlation function. The non-robust correlation shows its bias towards the outlying curves at parts of the domain where significant differences arise.
@@ 6.2 Fertility data
The second data set consists of age-specific fertility rates (ASFR) across different countries. ASFR represents the number of women giving birth per $1,000$ within specific age groups, providing a detailed view of how fertility changes with age. The resulting curves are analyzed as density data because we are interested in the underlying distribution across the different age groups rather than the magnitude of it. We examine data from the Human Fertility Database, available at https://www.humanfertility.org/Home/Index, focusing on the years $1956$, $1990$, and $2019$, to gain insight into the evolution of fertility dynamics.
In this example, we apply Algorithm 1 with 21 compositional splines as basis functions. The data is visualized in Figure 9. Outlying countries, according to their squared RDMD, are represented by dashed curves. During the early stages of this period, peak fertility was primarily achieved by young women in their early 20s. Large families were common, reflected in a rather slow decrease in the majority of the fertility curves. During the following decades, mostly influenced by the introduction of reliable contraceptives and a change in the mindset about life quality, later first pregnancies became more and more common. Throughout Europe, this phenomenon started earlier in the West than in the East, as can be seen by the fertility curves corresponding to the year $1990.$ Since then, the peak fertility has steadily increased and sits now at around 30 years. For more details, we refer to the insightful article by Beaujouan (2020) and references therein.
While the variability in the Bayes space seems to be present within the peaks, the upper row of Figure 9 shows that the main variability within the clr space is significantly higher in the tails of the distribution; see also Hron et al. (2016).
@@
In Figure 10, the first two principal functions are displayed. In the brackets, we included the corresponding percentage of explained variance. For each year, the first component mainly captures the variability in the tails of the densities. Here, the principal functions focus on the left tail and the following ages until the early 20s. The second principal function contains significantly less information, mainly explaining the right tail and the peak fertility.
@@
Following the construction of robust principal components, we look at the corresponding estimated principal component scores $\hat{s}_{j}$ associated with each country. In Figure 11, the first two robust scores for each country are visualized. For 1956, the outliers, as indicated by the triangles, all lay at the border of the data cloud. For Japan, the first and second scores are exceedingly high, indicating not only relatively low fertility in younger age groups but also proportionally very high peak fertility. For the outliers of 1990 and 2019, we obtain a similar picture dominated by mostly extreme scores. For 1990, this mainly reflects countries that have an unusually high or low fertility for younger and older age groups. Similarly, in 2019, these extreme values explain high fertility rates at younger ages as well as late peak fertility.
Figure 11: Robust PC scores for the first two eigenfunctions for the years $1956,1990,2019$. Circles correspond to regular observations, while the triangles indicate outlying countries.
@@ 7 Conclusions and discussion
Despite the attention that the Bayes space methodology for density data analysis has attracted over the last decades, little focus has been paid to robust frameworks necessary for meaningful analysis in the presence of anomalies. Therefore, this paper introduces robust density PCA (RDPCA) as a methodology to robustly estimate the covariance operator and functional principal components of densely sampled univariate density data. Within this newly proposed method, we extend the notion of a functional Mahalanobis distance (Berrendero et al., 2020; Galeano et al., 2015) to the Bayes space. The resulting regularized density Mahalanobis distance (RDMD) leads to a robust covariance estimation based on a subsample of those curves that best fit the underlying distribution. Based on this estimate, robust PCs for density function can be determined as summarized in Algorithm LABEL:alg:alg1. If the data can be expressed in a fixed basis, we provide an equivalent approach in Algorithm 1. During simulations, we demonstrated the efficiency of RDPCA as well as the advantages it has against a non-robust approach in the presence of outliers. Furthermore, if outliers arise within the tails of the densities, the clr transformation can amplify these deviations. Especially in these scenarios, we observed that common robust procedures that do not respect the nature of the data would fail.
@@
One of our initial assumptions was that the data had been observed at a dense grid. In the case where the data is only available at a sparse grid, further extensions would require smoothing by an appropriate basis, e.g., CB-splines (Machalová et al., 2021) specifically developed for Bayes spaces. Naturally, the next step from univariate data would be to consider multivariate densities. During multivariate functional data analysis, each observation contains the recording of several “functional" variables. In this setting, not only is the covariance between different time points considered, but also the relation between individual variables. If the second dimension is also continuously observed, the observations are so-called random surfaces. As these statistical fields show increasingly practical relevance (Berrendero et al. (2011); Górecki et al. (2018); Dai and Genton (2018); Masak and Panaretos (2023) and references therein), expanding RDPCA to these areas seems worthwhile. Calculating the RDMD involved regularizing by a suitable operator that smooths out unwanted noise components while keeping the relevant signal within the (uncontaminated) data unaffected. Another class of meaningful operators that utilize the functional nature of the data are differential operators, often used in Tikhonov regularization. One has to keep in mind that these operations have yet to be defined for Bayes spaces. Next to PCA, the regularized Mahalanobis distance could be used for concepts like linear or quadratic discriminant analysis for the classification of density data. As these methods rely on similarity measures involving several covariance operators, the robust classification of densities based on the RDMD would be suitable.
@@
Supplement: The supplement contains auxiliary results, further developing ideas on RDPCA. Next, it provides additional simulation analysis corresponding to Section 5.1 and 5.2, and a further analysis of the fertility data from Section 6.2. The supplement finishes with the proofs of theoretical results.